### v1

@@ 5 Natural revision in topos quantum logic
For any $\overline{S}\in H$, we have $\overline{S}\subseteq L$ and $1=L$. Hence, $1\wedge\overline{S}=\overline{S}\wedge 1=\overline{S}$ holds.
@@
For any $\overline{S},\overline{T}\in H$ and any $V\in\mathcal{V(\mathcal{H})}$, $(\overline{S}\wedge\overline{T})_{V}=\overline{S}_{V}\cap\overline{T}_{V}\subseteq\overline{S}_{V}$. So $\overline{S}\wedge\overline{T}\leq\overline{S}$ holds.
@@
This revision operation takes the elements within a propositional lattice as operators and performs the $\wedge$ operation of Heyting algebra on new information and the antecedents of the original consequence relation in the lattice to obtain new antecedents within the same lattice. These new antecedents correspond to a new consequence relation.
Lemma 2.
The $\wedge_{s}$ operation outside a Heyting algebra can be viewed as another type of revision operation, referred to as operator-level revision. Every formula corresponding to a closed space $a\in L$ induces an operator on $\mathcal{C}$
@@
$\bar{\varphi^{*}_{a}}(\vdash_{H_{1}})=\vdash_{H_{2}}\quad\Leftrightarrow\quad a\wedge_{s}A_{H_{1}}=A_{H_{2}}$
in which $A_{H_{1}}$ and $A_{H_{2}}$ represents antecedents of $\vdash_{H_{1}}$ and $\vdash_{H_{2}}$ respectively.
@@
This revision operation takes the elements in a property lattice as operators and performs the Sasaki projection $\varphi^{*}_{a}$ on new information and the antecedents of the original consequence relation in a propositional lattice to obtain new antecedents within another propositional lattice. These new antecedents also correspond to a new consequence relation.
The elements in the two propositional lattices are identical, but the truth-value assignments of the elements may differ.
Theorem 3.
Given a consequence relation $\vdash$ within a propositional lattice $H$. Let $\alpha\in H$ and $a\in L$ be formulas. There are two types of revision operators, called object-level revision and operator-level revision respectively, to revise the consequence relation. The former, denoted by $\bar{\alpha}$, maps the consequence relation to another one within the same lattice $H$. The latter, denoted by $\bar{\varphi^{*}_{a}}$, maps the consequence relation to another one within a different propositional lattice.
@@
When we use formulas as operators to revise the consequence relation, there are two sources of formulas: those from within the propositional lattice and those from outside. When a formula from within the propositional lattice is used as an operator to revise the consequence relation, the revision operator is applied to the antecedents of the consequence relation based on the $\wedge$ operation within Heyting algebra. This results in a new consequence relation within the same lattice. When a formula from outside the propositional lattice is used as an operator, the selected formula corresponds to a closed subspace, namely a projection operation, and is used as a parameter with Sasaki projection $\varphi_{a}^{*}$ operation to revise the antecedents of the original consequence relation. This results in a new consequence relation within another propositional lattice.
@@
In a propositional lattice, there exists an element $C$ that is assigned to true, which corresponds to a consequence relation $\vdash_{C}$ based on an antecedent $\{C\}$. For the former equation, upon adding new information “element $B$ is assigned to true”, an object-level revision operation $\wedge$ is performed on consequence relation $\vdash_{C}$ to obtain antecedents $\{B,C\}$, which correspond to consequence relation $\vdash_{B\wedge C}$. Then, upon adding new information “projecting onto space $a$”, an operator-level revision operation $\wedge_{s}$ is employed to revise consequence relation $\vdash_{B\wedge C}$, resulting in an antecedent $\{\varphi^{*}_{a}(B\wedge C)=D\}$ and corresponding to consequence relation $\vdash_{D}$. For the latter, firstly upon adding new information “projecting onto space $a$”, an operator-level revision operation $\wedge_{s}$ is employed to revise consequence relation $\vdash_{C}$, resulting in an antecedent $\{\varphi^{*}_{a}(C)\}$ and corresponding to consequence relation $\vdash_{\varphi^{*}_{a}(C)}$. Then, upon adding new information “element $B$ is assigned to true”, an object-level revision operation $\wedge$ is performed on consequence relation $\vdash_{\varphi^{*}_{a}(C)}$ to obtain an antecedent $\{B\wedge\varphi^{*}_{a}(C)=E\}$, which corresponds to consequence relation $\vdash_{E}$. As shown in the Figure 1.
Figure 1: Natural revision
@@
Both object-level revision and operator-level revision involve changing the truth values of elements in the propositional lattice. The difference lies in that object-level revision directly adds a new element assigned to be true, and performs $\wedge$ operation on the original elements to change their truth values. On the other side, although it is known from $a\wedge_{s}b\leq a$ that the new element $a$ is assigned to true, for operator-level revision, the new element is applied to elements of the original lattice by $\wedge_{s}$ operation to obtain elements of another lattice which is assigned to true.
@@
If we regard natural revision as a consequence revision system (CRS), this system should have two revision operations, $\wedge$ and $\wedge_{s}$, which correspond to two internalizing connectives, $\rightarrow$ and $\rightsquigarrow_{s}$ respectively. However, for the $a\wedge_{s}\_$ operation, since the element $a$ is an external operator and should be a parameter of this operation, it is denoted as $\varphi_{a}^{*}(\_)$. Correspondingly, the $a$ in the internalizing connective $a\rightsquigarrow_{s}\_$ should also be a parameter, denoted as $\varphi_{a,*}(\_)$. In this manner, the revision operator and so-called internalizing connective obtained by Sasaki adjoint are not suitable for the formalism of internalizing connectives. That is, the modus ponens obtained by Sasaki adjoint in standard quantum logic is not suitable for ordinary truth-value reasoning. The consequence relation only occurs within the propositional lattice, namely the formula set $Fml$ in $\vdash\subseteq Fml\times Fml$ needs to be selected from the elements in the same propositional lattice. The propositional lattice has the $\wedge$ operation of Heyting algebra, and new information can transform the formula set $Fml_{1}$ into another formula set $Fml_{2}$ in the same lattice. On the other hand, the Sasaki projection operation is an external operation of the propositional lattice, and new information can transform one formula set $Fml_{H_{1}}$ in the lattice $H_{1}$ into another formula set $Fml_{H_{2}}$ in another lattice $H_{2}$.
In summary, truth-value reasoning in quantum logic occurs both within a single propositional lattice and between different propositional lattices. The internal truth-value reasoning within a single lattice is referred to as static quantum logic, while the truth-value reasoning between two propositional lattices is referred to as dynamic quantum logic. Consequence relations can be revised using two revision operators: object-level revision within a propositional lattice and operator-level revision between propositional lattices. The former uses the $\wedge$ operation of Heyting algebra, while the latter uses the Sasaki projection $\varphi^{*}_{a}$ operation of Sasaki adjoint. By employing these revision operations, we can add new information to the antecedents of a consequence relation and revise it accordingly, leading to new consequence relations within the same or different propositional lattices.
@@ Conclusion
In this paper, we point out that the revision theory based on standard quantum logic cannot correspond to truth-value reasoning in quantum logic. The reasoning rules generated by Sasaki adjoint in standard quantum logic are not truth-value reasoning within the system, but truth-value reasoning obtained by taking external elements as parameters of the operation. To accurately describe truth-value reasoning in quantum logic, it is necessary to reconsider truth-value assignment in quantum logic. In this paper, we use topos quantum theory to characterize the truth-value assignment of quantum logic and indicate that the truth-value assignment proposed by Isham et al. in topos quantum theory does not apply to truth-value reasoning in the Heyting algebra sense. To give truth-value reasoning that satisfies Heyting algebra, we redefine truth-value assignment in topos quantum theory by taking the downset of a property as its inner daseinisation in all contexts, construct a Heyting algebra with downsets as elements, and assign truth values to downsets satisfies the law of non-contradiction. In this formalism, truth-value reasoning in the sense of Heyting algebra is carried out. We call such Heyting algebra a propositional lattice. Considering Sasaki adjoint in the sense of the propositional lattice, we regard the Sasaki adjoint as an operation that selects a property from the property lattice as a parameter and maps elements in one propositional lattice to those in another propositional lattice. By using internal and external operations in Heyting algebra, the reasoning rules of quantum logic can be characterized. Quantum logic is determined by both static and dynamic reasoning.
@@
After giving truth-value assignment and reasoning rules for quantum logic, we consider giving a natural revision theory based on quantum logic. We correspond the static and dynamic reasoning in quantum logic to two revision operations, which are object-level revision and operator-level revision, respectively. The static reasoning corresponds to object-level revision, which is a revision based on the $\wedge$ operation inside the Heyting algebra. The dynamic reasoning corresponds to operator-level revision, which is a revision based on the $\wedge_{s}$ operation outside the Heyting algebra. The consequence relation $\vdash\subseteq Fml\times Fml$ arises within a propositional lattice and the set of formulas $Fml$ is constituted by elements in the propositional lattice. Object-level revision revises the consequence relation to another consequence relation in the same propositional lattice, while operator-level revision revises the consequence relation within one propositional lattice to another consequence relation in another propositional lattice with the same elements but different truth-value assignments.
@@
The study of truth-value assignment and truth-value reasoning in intuitionistic quantum logic can be extended to mixed states of quantum mechanics. The investigation of how entanglement and superposition phenomena in quantum mechanics manifest in quantum logic is worth studying. In quantum computation, unitary operators represent evolution operations and can also serve as dynamic operators. This further expands the natural revision theory and provides a theoretical foundation for quantum computation.

### v2

@@ 5. Natural revision in topos quantum logic
For any $\overline{S}\in D(L)$, we have $\overline{S}\subseteq L$ and $1=L$. Hence, $1\wedge\overline{S}=\overline{S}\wedge 1=\overline{S}$ holds.
@@
For any $\overline{S},\overline{T}\in D(L)$ and any $V\in\mathcal{V(\mathcal{H})}$, $(\overline{S}\wedge\overline{T})_{V}=\overline{S}_{V}\cap\overline{T}_{V}\subseteq\overline{S}_{V}$. So $\overline{S}\wedge\overline{T}\leq\overline{S}$ holds.
@@
This revision operation takes the elements within Heyting algebra $D(L)$ as operators and performs the $\wedge$ operation of Heyting algebra on new message and the antecedents of the original consequence relation to obtain new antecedents within the same lattice. These new antecedents correspond to a new consequence relation.
Lemma 5.2.
The $\wedge_{s}$ operation outside a Heyting algebra can be viewed as another type of revision operation, referred to as dynamic revision. Every formula corresponding to a closed space $a\in L$ induces an operator on $\mathcal{C}$
@@
$\bar{\varphi^{*}_{a}}(\vdash_{D(L)_{1}})=\vdash_{D(L)_{2}}\quad\Leftrightarrow\quad a\wedge_{s}A_{D(L)_{1}}=A_{D(L)_{2}}$
in which $A_{D(L)_{1}}$ and $A_{D(L)_{2}}$ represents antecedents of $\vdash_{D(L)_{1}}$ and $\vdash_{D(L)_{2}}$ respectively.
@@
This revision operation takes a element $a$ in a property lattice as operator and performs the Sasaki projection $\varphi^{*}_{a}$ on new message and the antecedents of the original consequence relation in a lattice to obtain new antecedents within another lattice. These new antecedents also correspond to a new consequence relation.
It is worth noting that elements in the two lattices are identical, but the truth-value assignments of the elements may differ.
Theorem 5.3.
Given a consequence relation $\vdash$ within a Heyting algebra $H$. Let $\alpha\in H$ and $a\in L$ be formulas. There are two types of revision operators, called static and dynamic revision respectively, to revise the consequence relation. The former, denoted by $\bar{\alpha}$, maps the consequence relation to another one within the same lattice $H$. The latter, denoted by $\overline{\varphi^{*}_{a}}$, maps the consequence relation to another one within a different lattice.
@@
When we use formulas as operators to revise the consequence relation, there are two sources of formulas: those from within the Heyting algebra and those from outside. When a formula from within the Heyting algebra is used as an operator to revise the consequence relation, the revision operator is applied to the antecedents of the consequence relation based on the $\wedge$ operation within Heyting algebra. This results in a new consequence relation within the same lattice. When a formula from outside the Heyting algebra is used as an operator, the selected formula corresponds to a closed subspace, namely a projection operation, and is used as a parameter $a$ with Sasaki projection $\varphi_{a}^{*}$ operation to revise the antecedents of the original consequence relation. This results in a new consequence relation within another lattice.
@@
In a Heyting algebra, there exists an element $C$ that is assigned to true, which corresponds to a consequence relation $\vdash_{C}$ based on an antecedent $\{C\}$. For the former equation, upon adding new message “element $B$ is assigned to true”, a static revision operation $\wedge$ is performed on consequence relation $\vdash_{C}$ to obtain antecedents $\{B,C\}$, which correspond to consequence relation $\vdash_{B\wedge C}$. Then, upon adding new message “projecting onto space $a$”, a dynamic revision operation $\wedge_{s}$ is employed to revise consequence relation $\vdash_{B\wedge C}$, resulting in an antecedent $\{\varphi^{*}_{a}(B\wedge C)=D\}$ and corresponding to consequence relation $\vdash_{D}$. For the latter, firstly upon adding new message “projecting onto space $a$”, a dynamic revision operation $\wedge_{s}$ is employed to revise consequence relation $\vdash_{C}$, resulting in an antecedent $\{\varphi^{*}_{a}(C)\}$ and corresponding to consequence relation $\vdash_{\varphi^{*}_{a}(C)}$. Then, upon adding new message “element $B$ is assigned to true”, a static revision operation $\wedge$ is performed on consequence relation $\vdash_{\varphi^{*}_{a}(C)}$ to obtain an antecedent $\{B\wedge\varphi^{*}_{a}(C)=E\}$, which corresponds to consequence relation $\vdash_{E}$. As shown in the Figure 1.
Figure 1. Natural revision
@@
Both static and dynamic revision involve changing the truth-value assignments of elements in the Heyting algebra. The difference lies in that static revision directly adds a new element assigned to be true, and performs $\wedge$ operation on the original elements to change their truth values. On the other side, although it is known from $a\wedge_{s}b\leq a$ that the new element $a$ is assigned to true, for dynamic revision, the new element is applied to elements of the original lattice by $\wedge_{s}$ operation to obtain elements of another lattice which is assigned to true.
@@
If we regard natural revision as a consequence revision system (CRS), this system should have two revision operations, $\wedge$ and $\wedge_{s}$, which correspond to two internalizing connectives, $\rightarrow$ and $\rightsquigarrow_{s}$ respectively. However, for the $a\wedge_{s}\_$ operation, since the element $a$ is a parameter of external operation, it is denoted as $\varphi_{a}^{*}(\_)$. Correspondingly, the $a$ in the internalizing connective $a\rightsquigarrow_{s}\_$ should also be a parameter, denoted as $\varphi_{a,*}(\_)$. In this manner, the revision operator and so-called internalizing connective obtained by Sasaki adjoint are not suitable for the formalism of internalizing connectives. That is, the modus ponens obtained by Sasaki adjoint in standard quantum logic is not suitable for ordinary truth-value reasoning. The consequence relation only occurs within the Heyting algebra, namely the formula set $Fml$ in $\vdash\subseteq Fml\times Fml$ needs to be selected from the elements in the same lattice. The Heyting algebra has the $\wedge$ operation, and new message can transform the formula set $Fml_{1}$ into another formula set $Fml_{2}$ in the same lattice. On the other hand, the Sasaki projection is an external operation of the Heyting algebra, and new message can transform one formula set $Fml_{H_{1}}$ in the lattice $H_{1}$ into another formula set $Fml_{H_{2}}$ in another lattice $H_{2}$.
@@ Conclusion
In this paper, we point out that the revision theory based on standard quantum logic cannot correspond to truth-value reasoning in quantum logic. The reasoning rules generated by Sasaki adjoint in standard quantum logic are not truth-value reasoning within the system, but truth-value changes obtained by taking external elements as parameters of the operation. To accurately describe truth-value reasoning in quantum logic, it is necessary to reconsider truth-value assignment in quantum logic. In this paper, we use topos quantum theory and Heyting algebra to characterize the truth-value reasoning of quantum logic and indicate that the methodology proposed by Isham et al. in topos quantum theory does not apply to truth-value reasoning in the Heyting algebra sense. To give truth-value reasoning that satisfies Heyting algebra, we redefine the structure in topos quantum theory by taking the downset of a property as its inner daseinisation in all contexts, construct a Heyting algebra with downsets as elements, and assign truth values to downsets satisfies the law of non-contradiction. In this formalism, truth-value reasoning in the sense of Heyting algebra is carried out. Moreover, we regard the Sasaki adjoint as an operation that selects a property from the property lattice as a parameter and maps elements in one Heyting algebra to those in another. By using internal and external operations in Heyting algebra, the reasoning rules of quantum logic can be characterized. Quantum logic is determined by both static and dynamic reasoning.
@@
After giving truth-value assignment and reasoning rules for quantum logic, we consider giving a natural revision theory based on quantum logic. We correspond the static and dynamic reasoning in quantum logic to two revision operations, which are static and dynamic revision, respectively. The static reasoning corresponds to a revision based on the $\wedge$ operation inside the Heyting algebra. The dynamic reasoning corresponds to a revision based on the $\wedge_{s}$ operation outside the Heyting algebra. The consequence relation $\vdash\subseteq Fml\times Fml$ arises within a Heyting algebra and the set of formulas $Fml$ is constituted by elements in the lattice. Static revision revises the consequence relation to another consequence relation in the same lattice, while dynamic revision revises the consequence relation within one lattice to another consequence relation in another lattice with the same elements but different truth-value assignments.
@@
The research of truth-value assignment and truth-value reasoning in intuitionistic quantum logic can be extended to mixed states of quantum mechanics. The investigation of how entanglement and superposition phenomena in quantum mechanics manifest in quantum logic is worth studying. In quantum computation, unitary operators represent evolution operations and can also serve as dynamic operators. This further expands the natural revision theory and provides a theoretical foundation for quantum computation.